OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3. Choices y - 6 = -3(x + 1) 3x + y = 3 y = -3x + 3 y - 6 = -3x - 3

OpenStudy (anonymous):

a

OpenStudy (anonymous):

y-y1=a(x-x1) a is slope and put -1 as x that give 6 as y

OpenStudy (ash2326):

Standard equation of a line passing through a point and with slope m (y-y1)=m(x-x1) her the point is -1,6 and slope -3 (y-6)=-3(x+1) option A